### v2

# Deep SPI: Safe Policy Improvement via World Models

Abstract
Safe policy improvement (SPI) offers theoretical control over policy updates, yet existing guarantees largely concern offline, tabular reinforcement learning (RL). We study SPI in general online settings, when combined with world model and representation learning. We develop a theoretical framework showing that restricting policy updates to a well-defined neighborhood of the current policy ensures monotonic improvement and convergence. This analysis links transition and reward prediction losses to representation quality, yielding online, “deep” analogues of classical SPI theorems from the offline RL literature. Building on these results, we introduce DeepSPI, a principled on-policy algorithm that couples local transition and reward losses with regularised policy updates. On the ALE-57 benchmark, DeepSPI matches or exceeds strong behaviorals, including PPO and DeepMDPs, while retaining theoretical guarantees.
1 Introduction
Reinforcement learning (RL) trains agents to act in complex environments through trial and error (Sutton and Barto, 2018). To scale to high-dimensional domains, modern approaches rely on function approximation, making representation learning (Echchahed and Castro, 2025) essential for constructing latent spaces where behaviorally similar states are mapped close together and policies and value functions become easier to estimate. A complementary approach is model learning, where a predictive model of the environment is trained (Ha and Schmidhuber, 2018). Such models can be leveraged for planning, generating simulated experience, or improving value estimates (Hafner et al., 2021; Schrittwieser et al., 2020; Xiao et al., 2019).
In the online setting, where the agent updates its policy during interaction, avoiding catastrophic errors is critical. Two key challenges arise: out-of-trajectory (OOT) world models and confounding policy updates. OOT issues arise when the world model fails to capture rarely visited regions of the state space, leading to unreliable predictions and unsafe updates when the latent policy explores these regions (Suau et al., 2024). Confounding updates occur when both the policy and its underlying representation are updated simultaneously: poor representations can lock the agent into suboptimal behavior, while the policy itself prevents corrective updates to the representation. Safe Policy Improvement (SPI) mitigates such risks by ensuring that new policies are not substantially worse than their predecessors (Thomas et al., 2015). Classical SPI methods provide rigorous results in tabular MDPs but depend on exhaustive state–action coverage, making them unsuitable for continuous or high-dimensional spaces.
We address this gap by directly connecting representation and model learning with safe policy improvement in complex environments with general state spaces. Our contributions are threefold. First, we introduce a novel neighborhood operator that constrains policy updates, enabling policy improvement with convergence guarantees. Second, we combine this operator with principled model losses to bound the gap between a policy’s performance in the world model and in the true environment, thereby enabling safe policy improvement in complex MDPs. This analysis also shows that our scheme enforces representation quality by ensuring that states with similar values remain close in the learned latent space. Third, we connect our theory to PPO (Schulman et al., 2017) and propose DeepSPI, a practical algorithm that achieves strong empirical performance on the Arcade Learning Environment (ALE; Bellemare et al. 2013) while retaining theoretical guarantees.
1.1 Related Work
Regularizing policy improvements. Regularized updates, as in TRPO, PPO, and related analyses, are now standard for stabilizing policy optimization (Schulman et al., 2015; 2017; Geist et al., 2019; Kuba et al., 2022). Our work extends this perspective to the joint training of a world model and a representation, where we constrain policy updates in a principled neighborhood while controlling model quality through transition and reward losses.
SPI methods provide principled guarantees on policy updates from fixed datasets (offline RL) (Thomas et al., 2015; Ghavamzadeh et al., 2016a; Laroche et al., 2019; Simão et al., 2020; Castellini et al., 2023). These methods assume tabular state spaces and offline data, where error bounds must hold globally across all state–action pairs, often via robust MDP formulations (Iyengar, 2005; Nilim and Ghaoui, 2005). Our setting is fundamentally different: we study online RL with high-dimensional inputs, where such global constraints are intractable. We take inspiration from the SPI literature but introduce local, on-policy losses that make safe improvement feasible in practice. In spirit, other model-based methods share the goal of providing SPI-like guarantees in more general settings, but are again purely offline, omit any form of representation learning, and rely on assumptions that differ substantially from ours (Yu et al., 2020; 2021; Kidambi et al., 2020).
Representation learning and model-based RL. Auxiliary transition and reward prediction losses are central to many model-based methods, from DeepMDP to Dreamer and related world-model approaches (Gelada et al., 2019; Hafner et al., 2021). In particular, the losses we consider for learning transitions and rewards generalize a wide range of objectives used across the model-based RL literature (François-Lavet et al., 2019; van der Pol et al., 2020; Kidambi et al., 2020; Delgrange et al., 2022; Dong et al., 2023; Alegre et al., 2023). Conceptually, our representation guarantees are closely connected to classical notions of state abstraction in MDPs (Li et al., 2006) and to bisimulation (Larsen and Skou, 1991; Desharnais et al., 1998). Building on bisimulation, prior work develops representations that group states into areas in which the agent is guaranteed to behave similarly under the current policy (Castro, 2020; Zhang et al., 2021; Castro et al., 2021; Agarwal et al., 2021a; Avalos et al., 2024). By contrast, we directly link representation quality and model accuracy to our safe policy improvement analysis, yielding tractable guarantees in the online setting.
2 Background
In the following, given a measurable space $\mathcal{X}$, we write $\Delta\mathopen{}\mathclose{{\left(\mathcal{X}}}\right)$ for the set of distributions over $\mathcal{X}$. For any distribution $\mu\in\Delta\mathopen{}\mathclose{{\left(\mathcal{X}}}\right)$, we denote by $\text{supp}\mathopen{}\mathclose{{\left(\mu}}\right)$ its support.
Markov Decision Processes (MDPs) offer a formalism for sequential decision-making under uncertainty. Formally, an MDP is a tuple of the form $\mathcal{M}=\langle\mathcal{S},\mathcal{A},P,{R},s_{\mathit{I}},\gamma\rangle$ consisting of a set of states $\mathcal{S}$, actions $\mathcal{A}$, a transition function $P:\mathcal{S}\times\mathcal{A}\to\Delta\mathopen{}\mathclose{{\left(\mathcal{S}}}\right)$, a bounded reward function ${R}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ with $\left\|{R}\right\|_{\infty}={R}_{\textsc{max}}$, an initial state $s_{\mathit{I}}\in\mathcal{S}$, and a discount factor $\gamma\in[0,1)$. Unless otherwise stated, we generally assume that $\mathcal{S}$ and $\mathcal{A}$ are compact. An agent interacting in $\mathcal{M}$ produces trajectories, i.e., infinite sequences of states and actions $\mathopen{}\mathclose{{\left(s_{t},a_{t}}}\right)_{t\geq 0}$ visited along the interaction so that $s_{0}=s_{\mathit{I}}$ and $s_{t+1}\sim P\mathopen{}\mathclose{{\left({\cdotp}\mid s_{t},a_{t}}}\right)$ for all $t\geq 0$.
At each time step $t$, the agent selects an action according to a (stationary) policy $\pi:\mathcal{S}\to\Delta\mathopen{}\mathclose{{\left(\mathcal{A}}}\right)$ mapping states to distributions over actions. Running an MDP under $\pi$ induces a unique probability measure $\mathbb{P}_{\pi}$ over trajectories (Revuz, 1984), with associated expectation operator $\mathbb{E}_{\pi}$; we write $\mathbb{E}_{\pi}[\cdot\mid s_{0}=s]$ when the initial state is fixed to $s\in\mathcal{S}$. A policy has full support if $\text{supp}\mathopen{}\mathclose{{\left(\pi(\cdot\mid s)}}\right)=\mathcal{A}$ for all $s\in\mathcal{S}$, and we denote the set of all policies by $\Pi$. A stationary measure of $\pi$ is a distribution over states visited under $\pi$, and is defined as a solution of $\xi_{\pi}(\cdot)=\mathbb{E}_{s\sim\xi_{\pi}}\mathbb{E}_{a\sim\pi(\cdot\mid s)}[P(\cdot\mid s,a)]$. Such a measure is often assumed to exist in continual RL (Sutton and Barto, 2018), is unique in episodic RL (Huang, 2020), and defines the occupancy measure in discounted RL (Metelli et al., 2023). [^1] Details on the formalization of episodic processes and value functions can be found in Appendix A.
Value functions. The performance of the agent executing a policy $\pi\in\Pi$ in each single state $s\in\mathcal{S}$ can be evaluated through the value function $V^{\pi}\mathopen{}\mathclose{{\left(s}}\right)=\mathop{\mathbb{E}}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}{R}(s_{t},a_{t})\mid s_{0}=s\right]$. The goal of an agent is to maximize the return from the initial state, given by $\rho\mathopen{}\mathclose{{\left(\pi,\mathcal{M}}}\right)=V^{\pi}\mathopen{}\mathclose{{\left(s_{\mathit{I}}}}\right)$. To evaluate the quality of any action $a\in\mathcal{A}$, we consider the action value function $Q^{\pi}\mathopen{}\mathclose{{\left(s,a}}\right)={R}(s,a)+\gamma\mathop{\mathbb{E}}_{s^{\prime}\sim P(\cdot\mid s,a)}{{V}^{\pi}\mathopen{}\mathclose{{\left(s^{\prime}}}\right)}$, being the unique solution of Bellman’s equation with ${V}^{\pi}\mathopen{}\mathclose{{\left(s}}\right)=\mathop{\mathbb{E}}_{a\sim\pi\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)}{Q^{\pi}\mathopen{}\mathclose{{\left(s,a}}\right)}$. Alternatively, any given action can be evaluated through the advantage function $A^{\pi}\mathopen{}\mathclose{{\left(s,a}}\right)=Q^{\pi}\mathopen{}\mathclose{{\left(s,a}}\right)-V^{\pi}\mathopen{}\mathclose{{\left(s}}\right)$, giving the advantage of selecting an action over the current policy.
Representation learning in RL. In realistic environments, the state–action space is too large for tabular policies or value functions. Instead, deep RL employs an encoder $\phi\colon\mathcal{S}\to\,\overline{\!{\mathcal{\mathcal{S}}}}$ that maps states to a tractable latent space $\overline{\!{\mathcal{\mathcal{S}}}}$, from which value functions can be approximated. Learning such encoders is referred to as representation learning (Echchahed and Castro, 2025). To improve representations, agents are often trained with additional objectives, commonly auxiliary tasks requiring predictive signals. Policy-based methods then optimize a latent policy $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\colon\,\overline{\!{\mathcal{\mathcal{S}}}}\to\Delta\mathopen{}\mathclose{{\left(\mathcal{A}}}\right)$ jointly with $\phi$, executed in the environment as $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu(\cdot\mid\phi(s))$. By convention, we write $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu(\cdot\mid s)$ for $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\circ\phi(s)$ when $\phi$ is clear, and denote the set of all latent policies by $\mkern 1.5mu\overline{\mkern-1.5mu\Pi\mkern-1.5mu}\mkern 1.5mu$. For any $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\in\mkern 1.5mu\overline{\mkern-1.5mu\Pi\mkern-1.5mu}\mkern 1.5mu$, the composed policy $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\circ\phi$ belongs to $\Pi$.
Model-based RL augments policy learning with a world model $\overline{\!{\mathcal{M}}}=\langle{\,\overline{\!{\mathcal{\mathcal{S}}}},\mathcal{A},\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu,\,\overline{\!{{R}}},\bar{s}_{I},\gamma\rangle}$, which can improve (i) sample efficiency by generating trajectories (e.g., Hafner et al. 2021), (ii) value estimation through planning (e.g., Buckman et al. 2018), and (iii) representation learning by grouping states with similar behavior (e.g., Gelada et al. 2019; Zhang et al. 2021). When $\overline{\!{\mathcal{\mathcal{S}}}}=\mathcal{S}$, the model must replicate environment dynamics, which is often intractable. Instead, we focus on $\overline{\!{\mathcal{\mathcal{S}}}}$ defined by the learned representation $\phi$, so that $\overline{\!{\mathcal{M}}}$ becomes an abstraction of $\mathcal{M}$. Learning transition and reward functions then additionally serves as an auxiliary signal for the representation, encouraging states with similar behavior to map close in $\overline{\!{\mathcal{\mathcal{S}}}}$. Since $\overline{\!{\mathcal{\mathcal{S}}}}$ is the latent space, $\mkern 1.5mu\overline{\mkern-1.5mu\Pi\mkern-1.5mu}\mkern 1.5mu$ corresponds to the policies of $\overline{\!{\mathcal{M}}}$. We further assume $\overline{\!{\mathcal{\mathcal{S}}}}$ is equipped with a metric $\overline{\!{d}}\colon\,\overline{\!{\mathcal{\mathcal{S}}}}\times\,\overline{\!{\mathcal{\mathcal{S}}}}\to[0,\infty)$ to measure distances.
3 No way Home: when world models and policies go out of trajectories
World models are usually learned toward minimizing a reward loss $L_{{R}}$ and/or transition loss $L_{P}$ from experiences $\eta$ collected along the agent’s trajectories. Those experiences are either gathered in the form of a batch or a replay buffer $\mathcal{B}$. In general, the loss functions take the following form: $L_{{R}}=\mathop{\mathbb{E}}_{\eta\sim\mathcal{B}}f_{{R}}\mathopen{}\mathclose{{\left(\phi,\,\overline{\!{{R}}};\,\eta}}\right)$ and $L_{P}=\mathop{\mathbb{E}}_{\eta\sim\mathcal{B}}f_{P}\mathopen{}\mathclose{{\left(\phi,\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu;\,\eta}}\right)$, where $f_{R}$ (resp. $f_{P}$) assign a “cost” relative to the error between R{R} and $\overline{\!{{R}}}$ (resp. $P$ and $\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu$) according to the experiences $\eta$ and their representation. Henceforth, we refer to the policy $\pi_{\text{b}}$ used to insert experiences in $\mathcal{B}$ as the behavioral policy.
3.1 Out-of-trajectory world model
One may consider leveraging the model $\overline{\!{\mathcal{M}}}$ to improve the policy $\pi_{\text{b}}$. This can be achieved by directly planning a new policy $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$ in $\overline{\!{\mathcal{M}}}$ or drawing imagined trajectories in the world model to evaluate new actions and improve on sample complexity during RL. However, since the world model is learned from experiences stored in $\mathcal{B}$, we can only be certain of its average accuracy according to this data. This is problematic because some regions of the state space of $\mathcal{M}$ may have been rarely, or not at all, visited under $\pi_{\text{b}}$. In that case, the predictions made in $\overline{\!{\mathcal{M}}}$ might cause the agent to “hallucinate” inaccurate trajectories in the latent space and spoil the policy improvement. This problem, known as the out-of-trajectory (OOT) issue (Suau et al., 2024), arises when a policy in $\overline{\!{\mathcal{M}}}$ deviates substantially from $\pi_{\text{b}}$, which can render the model unreliable.
(a) A large MDP $\mathcal{M}$ whose state space is divided in four regions $\mathcal{S}=\bigcup_{i=1}^{4}\mathcal{S}_{i}$.
(b) A simple world model $\overline{\!{\mathcal{M}}}$ whose state space is $\overline{\!{\mathcal{\mathcal{S}}}}=\left\{\bar{s}_{1},\bar{s}_{2},\bar{s}_{3},\bar{s}_{3}^{\prime},\bar{s}_{4}\right\}$.
Figure 1: In $\mathcal{M}$, continuously playing $a_{1}$ in states from $\mathcal{S}_{1}$ eventually leads the agent to the region $\mathcal{S}_{2}$, and playing $a_{3}$ in $\mathcal{S}_{2}$ eventually leads the agent to $\mathcal{S}_{4}$ where a reward of $1$ is incurred at each time step, whatever the action played. Playing $a_{2}$ in $\mathcal{S}_{1}$ leads the agent to the region $\mathcal{S}_{3}$, where all actions incur negative rewards. Here, $\phi\mathopen{}\mathclose{{\left(s}}\right)=\bar{s}_{i}$ for any $s\in\mathcal{S}_{i}$ and $i=\left\{1,2,4\right\}$. For $s\in\mathcal{S}_{3}$, we have either $\phi\mathopen{}\mathclose{{\left(s}}\right)=\bar{s}_{3}$ or $\phi\mathopen{}\mathclose{{\left(s}}\right)=\bar{s}^{\prime}_{3}$.
To illustrate this problem, consider the world model of Figure 1. Assume the model is trained by collecting trajectories produced by $\pi_{\text{b}}$ in $\mathcal{M}$ where $\pi_{\text{b}}\mathopen{}\mathclose{{\left(a_{2}\mid s}}\right)\leq\epsilon$ for all $s\in\mathcal{S}_{1}$, with $\epsilon>0$. For a sufficiently small $\epsilon$, the region $\mathcal{S}_{3}$ in the original environment would remain largely unexplored while having almost no impact on the losses $L_{{R}},L_{P}$. Therefore, the representation of states in $\mathcal{S}_{3}$ ($\bar{s}_{3}$ and $\bar{s}^{\prime}_{3}$) may turn completely inaccurate. Here, the model incorrectly assigns a reward of $20$ to $\bar{s}^{\prime}_{3}$, whereas the true reward is strictly negative. Consequently, the optimal policy in $\overline{\!{\mathcal{M}}}$ deterministically selects $a_{2}$ in $\bar{s}_{1}$. When executed in the original environment, this policy drives the agent to $\mathcal{S}_{3}$ thereby degrading the behavioral policy $\pi_{\text{b}}$.
3.2 Confounding policy update
Updating both the representation and the policy solely from experience collected under a behavioral policy can degrade performance rather than improve it. In the same spirit as policy confounding (Suau et al., 2024), we call this phenomenon confounding policy update. The MDP in Figure 2 illustrates the issue.
Figure 2: MDP where the probability of transitioning from $s_{1}$ to $s_{2}$ is $1-\epsilon$, for $0<\epsilon<\nicefrac{{1}}{{4}}$.
The agent maps the states $s_{2}$ and $s_{3}$ to the same latent state $\bar{s}$, i.e. $\phi(s)=\bar{s}$ iff $s\in\{s_{2},s_{3}\}$. States $s_{1}$ and $s_{4}$ each have their own latent state. We consider the behavioral policy $\pi_{\text{b}}:=\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}\circ\phi$, where $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}$ is a stochastic policy with a small exploration rate $\zeta$:
$\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}(a_{1}\mid\bar{s})=1-\zeta,\qquad\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}(a_{2}\mid\bar{s})=\zeta,$ (1)
for $0<\zeta\ll\epsilon$. A good representation would ideally group states from which the agent behaves similarly. Because trajectories that reach $s_{3}$ and pick $a_{2}$ are unlikely, the two states appear identical under $\pi_{\text{b}}$: $|V^{\pi_{\text{b}}}(s_{2})-V^{\pi_{\text{b}}}(s_{3})|\!\approx\!0$. Therefore, this justifies using $\phi$ as representation for $\pi_{\text{b}}$, because the values of $s_{2}$ and $s_{3}$ are nearly identical: the agent exhibits close behaviors under $\pi_{\text{b}}$ from those states.
Suppose exploration under $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}$ eventually discovers that playing $a_{2}$ in $\bar{s}$ sometimes yields the +2+2 reward. Based on exploration data, an RL agent might therefore be tempted to change the latent policy to $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu(a_{2}\mid\bar{s})=1$ without modifying the representation $\phi$. With the representation still grouping $s_{2}$ and $s_{3}$, the new policy would now deterministically pick $a_{2}$ in both concrete states. Whenever the agent actually reaches $s_{3}$, it would receive the large negative reward $\nicefrac{{-2}}{{\epsilon}}$, which turns the overall return (from $s_{1}$) negative, thus worse than under $\pi_{\text{b}}$ even though $a_{2}$ is indeed optimal in $s_{2}$.
A solution to this problem would have been to split the representation of $s_{2}$ and $s_{3}$ in two distinct latent states. In general, representation and policy learning must be coupled since any change in the policy that alters the distribution over states can invalidate a previously adequate representation. However, in this example, the agent has no incentive to do so based on the experiences collected under $\pi_{\text{b}}$. As we will show below, updating both the policy and the representation jointly should be handled carefully to ensure policy improvement.
Our goal is to establish sufficient conditions to guarantee safe policy improvement during the RL process, either based on world models, state representations, or both, thus alleviating OOT world model and confounding policy update issues. Notice that, in the examples, both problems occur when performing aggressive updates from $\pi_{\text{b}}$ to a new policy $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$ (the mode of the distributions drastically shifts). Intuitively, smooth updates indeed ensure to alleviate those issues: constraining the policy search to policies “close” to $\pi_{\text{b}}$ (i) prevents hallucinations in parts of the world model that have been underexplored; (ii) reduces the risk of significantly degrading the return when updating the policy. While the benefits of regularizing policy improvements have already been both theoretically and practically justified (e.g., Geist et al. 2019; Kuba et al. 2022), their implications when mixing model-based and representation learning in RL have been underexplored.
Roadmap.
To rigorously address the OOT and confounding-update issues, the next sections develop the theoretical foundations of our approach, showing how controlled policy updates, local model losses, and representation stability interact. We briefly summarize how the main results connect.
Our analysis combines neighborhood-restricted policy updates, model-quality bounds, and representation guarantees. Sect. 4 introduces the neighborhood operator defining a trust region around the behavioral policy; restricting updates to this region ensures monotonic improvement and convergence (Thm. 1). Sect. 5 then links the reward and transition losses to value discrepancies: Thm. 2 shows that, when these losses are small, and updates remain in the neighborhood, the world model stays accurate under the learned representation. Combining these ingredients yields our first SPI result (Thm. 3), guaranteeing that direct policy updates in the world model translate to improvement under controlled error. Finally, Thm. 4 shows that the same loss-based control stabilizes the encoder, ensuring that value-distinct states remain separated in the latent space.
4 Your friendly neighborhood policy
Motivated by the intuition that constraining policy updates can mitigate OOT and confounding policy issues, we consider measuring the update as the importance ratio (IR) of the policies. This measure provides guarantees for constraining policy and representation updates, and with an appropriate optimisation scheme, ensures both policy improvement and convergence. In Section 5, we will further show that properly constraining the IR allows for safe policy improvements in world models while providing representation guarantees.
Let $\pi,\pi^{\prime}\in\Pi$, the extremal importance ratios are defined as $D^{\operatorname*{ext}}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)=\operatorname*{ext}\left\{\nicefrac{{\pi^{\prime}\mathopen{}\mathclose{{\left(a\mid s}}\right)}}{{\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)}}\colon s\in\mathcal{S},a\in\text{supp}\mathopen{}\mathclose{{\left(\pi\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)}}\right)\right\},\,\text{where }\operatorname*{ext}\in\left\{\inf,\sup\right\}.$ We define a neighborhood operator [^2] There are clear similarities between the IR, our neighborhood operator, and the PPO loss function (Schulman et al., 2017). We discuss this connection in Section 6. based on the IR, $\mathcal{N}^{C}\colon\Pi\to 2^{\Pi}$ for some constant $1<C<2$, establishing a trust region for policies updates that constraints the IR between 2C2-C and $C$:
$\mathcal{N}^{C}\mathopen{}\mathclose{{\left(\pi}}\right)=\left\{\pi^{\prime}\in\Pi\,\bigg|\,\begin{array}[]{ll}2-C\leq D^{\inf}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)\leq D^{\sup}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)\leq C,\\[5.0pt] \text{and }\,\text{supp}\mathopen{}\mathclose{{\left(\pi\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)}}\right)=\text{supp}\mathopen{}\mathclose{{\left(\pi^{\prime}\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)}}\right)\quad\forall s\in\mathcal{S}\end{array}\right\}\quad\quad\forall\pi\in\Pi.$ (2)
A critical question is whether an agent that restricts its policy updates to a defined neighborhood is truly following a sound policy improvement scheme. The following theorem shows that it does and further guarantees convergence.
Theorem 1.
(Policy improvement and convergence guarantees) Assume $\mathcal{S}$ and $\mathcal{A}$ are finite spaces. Let $\pi_{0}\in\Pi$ be a policy with full support and $\mathopen{}\mathclose{{\left(\pi_{n}}}\right)_{n\geq 0}$ be a sequence of policy updates defined as
$\pi_{n+1}\coloneqq\operatorname*{arg\,sup}_{\pi^{\prime}\in\mathcal{N}^{C}\mathopen{}\mathclose{{\left(\pi_{n}}}\right)}\,\mathop{\mathbb{E}}_{s\sim\mu_{\pi_{n}}}\mathop{\mathbb{E}}_{a\sim\pi^{\prime}\mathopen{}\mathclose{{\left({\cdotp}\mid{s}}}\right)}{A^{\pi_{n}}\mathopen{}\mathclose{{\left(s,a}}\right)},$ (3)
where $\mu_{\pi_{n}}$ is a sampling distribution with $\text{supp}\mathopen{}\mathclose{{\left(\mu_{\pi_{n}}}}\right)=\mathcal{S}$ for each $n\geq 0$. Then, the value function $V^{\pi_{n}}$ is monotonically improving, converges to $V^{*}$, and so is the return $\rho\mathopen{}\mathclose{{\left(\pi_{n},\mathcal{M}}}\right)$.
The proof consists in showing the resulting policy update scheme is an instance of mirror learning (Kuba et al., 2022), which yields the guarantees. Notice that since $\pi_{0}$ has full support, all the subsequent policies $\pi_{n}$ have full support as well. To maintain the guarantees, considering a stationary measure $\xi_{\pi_{n}}$ as the sampling distribution is only possible when $\text{supp}\mathopen{}\mathclose{{\left(\xi_{\pi_{n}}}}\right)=\mathcal{S}$. Note that this is always the case in episodic tasks (as the policy itself has full support). This is more generally true in ergodic MDPs (Puterman, 1994).
5 With great world models comes great representation
This section explains how the neighborhood operator of Eq. 2 enables safe policy improvement during world-model planning and representation updates in complex environments. Standard SPI methods ignore representation learning and require exhaustive state–action coverage in $\mathcal{B}$ to obtain guarantees, making them unsuitable for general state-action spaces. Even in finite domains, bounding the count of each state–action pair does not scale. Laroche et al. (2019) proposed baseline bootstrapping for under-sampled pairs, but their approach remains impractical in large-scale settings despite conceptual similarities to our operator. Further discussion of SPI limitations is provided in Appendix D.
Learning an accurate world model. SPI typically relies on optimizing a policy with respect to a latent model learned from the data stored in $\mathcal{B}$. In contrast to previous methods, our approach scales to high-dimensional feature spaces by (i) learning a representation $\phi$ and (ii) considering local error measures as opposed to global measures across the whole state-action space. We formalize them as tractable loss functions. Their local nature makes them compliant with stochastic gradient descent methods. Formally, given a distribution $\mathcal{B}\in\Delta\mathopen{}\mathclose{{\left(\mathcal{S}\times\mathcal{A}}}\right)$, we define the reward loss $L_{{R}}^{\mathcal{B}}$ and the transition loss $L_{P}^{\mathcal{B}}$ as
$L_{{R}}^{\mathcal{B}}\coloneqq\mathop{\mathbb{E}}_{s,a\sim\mathcal{B}}\left|{R}\mathopen{}\mathclose{{\left(s,a}}\right)-\,\overline{\!{{R}}}\mathopen{}\mathclose{{\left(\bar{s},a}}\right)\right|,$ $L_{P}^{\mathcal{B}}\coloneqq\mathop{\mathbb{E}}_{s,a\sim\mathcal{B}}\mathcal{W}\left(\phi_{\sharp}P\mathopen{}\mathclose{{\left({\cdotp}\mid s,a}}\right),\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left({\cdotp}\mid\phi\mathopen{}\mathclose{{\left(s}}\right),a}}\right)\right)$ (4)
where $\phi_{\sharp}P$ is the pushforward measure of $P$ by $\phi$, and $\mathcal{W}$ the Wasserstein distance (Vaserstein, 1969). $\mathcal{W}$ between $\mu,\nu\in\Delta\mathopen{}\mathclose{{\left(\,\overline{\!{\mathcal{\mathcal{S}}}}}}\right)$ is defined as $\mathcal{W}\left(\mu,\nu\right)=\inf_{\lambda\in\Lambda\mathopen{}\mathclose{{\left(\mu,\nu}}\right)}\mathop{\mathbb{E}}_{\mathopen{}\mathclose{{\left(\bar{s},\bar{s}^{\prime}}}\right)\sim\lambda}\;\,\overline{\!{d}}\mathopen{}\mathclose{{\left(\bar{s},\bar{s}^{\prime}}}\right)$, where $\Lambda(\mu,\nu)$ is the set of all couplings of $\mu$ and $\nu$. While the Wasserstein operator may seem scary at first glance, it generalizes over transition losses that can be found in the literature (cf. Sect. 1.1). In particular, when the latent space is discrete, this distance boils down to the total variation distance. Another notable case is when the transition dynamics are deterministic, in which case the transition loss reduces to $L_{P}^{\mathcal{B}}=\mathop{\mathbb{E}}_{s,a,s^{\prime}\sim\mathcal{B}}\,\,\overline{\!{d}}\mathopen{}\mathclose{{\left(\phi\mathopen{}\mathclose{{\left(s^{\prime}}}\right),\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left(\phi\mathopen{}\mathclose{{\left(s}}\right),a}}\right)}}\right)$. Finally, in general, a tractable upper bound can be obtained as $L^{\mathcal{B}}_{P}\leq\mathop{\mathbb{E}}_{s,a,s^{\prime}\sim\mathcal{B}}\mathop{\mathbb{E}}_{\bar{s}^{\prime}\sim\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left({\cdotp}\mid\phi\mathopen{}\mathclose{{\left(s}}\right),a}}\right)}{\,\overline{\!{d}}\mathopen{}\mathclose{{\left(\phi\mathopen{}\mathclose{{\left(s^{\prime}}}\right),\bar{s}^{\prime}}}\right)}$ (proof in Appendix C).
Lipschitz constants. To provide the guarantees, for any particular policy $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\in\mkern 1.5mu\overline{\mkern-1.5mu\Pi\mkern-1.5mu}\mkern 1.5mu$, we assume the world model is equipped with Lipschitz constants $K^{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}_{\,\overline{\!{{R}}}}$, $K^{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}_{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}$ defined as follows: for all $\bar{s}_{1},\bar{s}_{2}\in\,\overline{\!{\mathcal{\mathcal{S}}}}$,
$\left|\mathop{\mathbb{E}}_{a_{1}\sim\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left({\cdotp}\mid\bar{s}_{1}}}\right)}\,\overline{\!{{R}}}\mathopen{}\mathclose{{\left(\bar{s}_{1},a_{1}}}\right)-\mathop{\mathbb{E}}_{a_{2}\sim\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left({\cdotp}\mid\bar{s}_{2}}}\right)}\,\overline{\!{{R}}}\mathopen{}\mathclose{{\left(\bar{s}_{2},a_{2}}}\right)\right|$ $\leq K^{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}_{\,\overline{\!{{R}}}}\cdot\,\overline{\!{d}}\mathopen{}\mathclose{{\left(\bar{s}_{1},\bar{s}_{2}}}\right),$
$\mathcal{W}\left(\mathop{\mathbb{E}}_{a_{1}\sim\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left({\cdotp}\mid\bar{s}_{1}}}\right)}\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left({\cdotp}\mid\bar{s}_{1},a_{1}}}\right),\mathop{\mathbb{E}}_{a_{2}\sim\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left({\cdotp}\mid\bar{s}_{2}}}\right)}\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left({\cdotp}\mid\bar{s}_{2},a_{2}}}\right)\right)$ $\leq K^{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}_{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}\cdot\,\overline{\!{d}}\mathopen{}\mathclose{{\left(\bar{s}_{1},\bar{s}_{2}}}\right).$
Intuitively, the Lipschitzness of the latent reward and transition functions guarantees that the latent space is well-structured, so that nearby latent states exhibit similar latent dynamics. Gelada et al. (2019) control those bounds by adding a gradient penalty term to the loss and enforce Lipschitzness (Gulrajani et al., 2017). One can also obtain constrained Lipchitz constants as a side effect by enforcing the metric $\overline{\!{d}}$ to match the bisimulation distance in the latent space (Zhang et al., 2021). Interestingly, when the latent space is discrete, Lipschitz constants can be trivially inferred since $K^{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}_{\,\overline{\!{{R}}}}={2{R}_{\textsc{max}}}$ and $K^{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}_{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}=1$ (Delgrange et al., 2022). Note also that as the spaces are assumed compact, restricting to continuous functions ensures Lipschitz continuity.
For the sake of presentation, we restrict our attention to the following assumption for Thms. 2 and 3:
Assumption 1.
We assume that the agent operates in the episodic RL setting, i.e., we consider the standard RL framework where the environment is eventually reset with probability one.
Our results extend to general settings where a stationary distribution is accessible (c.f. Remark LABEL:rmk:episodic-assumption).
World model quality. Before introducing our safe policy improvement theorem, we first show that the local losses effectively measure the world model’s quality with respect to the original environment. Namely, their difference in return obtained under any latent policy in a well-defined neighborhood is bounded by the local losses derived from the reference, behavioral policy’s state-action distribution. This is formalized in the following theorem.
Theorem 2.
Suppose $\gamma>\nicefrac{{1}}{{2}}$ and $K^{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}_{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}<\nicefrac{{1}}{{\gamma}}$. Let $C\in\mathopen{(}1,\nicefrac{{1}}{{\gamma}}\mathclose{)}$, $\pi_{\text{b}}\in\Pi$ be the base policy, $\mathopen{}\mathclose{{\left(\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\circ\phi}}\right)\in\mathcal{N}^{C}\mathopen{}\mathclose{{\left(\pi_{\text{b}}}}\right)$ where $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\in\mkern 1.5mu\overline{\mkern-1.5mu\Pi\mkern-1.5mu}\mkern 1.5mu$ is a latent policy and $\phi\colon\mathcal{S}\to\,\overline{\!{\mathcal{\mathcal{S}}}}$ a state representation. Then,
$\left|\rho\mathopen{}\mathclose{{\left(\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\circ\phi,\mathcal{M}}}\right)-\rho\mathopen{}\mathclose{{\left(\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu,\,\overline{\!{\mathcal{M}}}}}\right)\right|\leq\textsc{AEL}\mathopen{}\mathclose{{\left(\pi_{\text{b}}}}\right)\cdot\frac{\nicefrac{{L_{{R}}^{\xi_{\pi_{\text{b}}}}}}{{\gamma}}+K_{V}\cdot L_{P}^{\xi_{\pi_{\text{b}}}}}{{\nicefrac{{1}}{{D^{\sup}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi_{\text{b}},\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}}\right)}}{-}\gamma}},$
where $\textsc{AEL}\mathopen{}\mathclose{{\left(\pi_{\text{b}}}}\right)$ denotes the average episode length when $\mathcal{M}$ runs under $\pi_{\text{b}}$, $K_{V}=\nicefrac{{K^{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}_{\,\overline{\!{{R}}}}}}{{\mathopen{}\mathclose{{\left(1-\gamma K^{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}_{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}}}\right)}}$, and $L_{{R}}^{\xi_{\pi_{\text{b}}}},L_{P}^{\xi_{\pi_{\text{b}}}}$ are the local losses of Eq. 4 over the stationary distribution $\xi_{\pi_{\text{b}}}$ induced by $\pi_{\text{b}}$.
In simpler terms, if the deviation (supremum IR, or SIR for short) between the behavioral policy and any new policy $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$ stays strictly lower than $\nicefrac{{1}}{{\gamma}}$, the gap in return between the environment and the world model for this new policy can be bounded using data collected via $\pi_{\text{b}}$. Minimizing local losses from $\pi_{\text{b}}$’s data ensures that refining the representation $\phi$ for $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$ improves model quality: when these losses vanish, $\mathcal{M}$ and $\overline{\!{\mathcal{M}}}$ are almost surely equivalent under $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$. The bound depends on the Average Episode Length (AEL), but even a loose upper bound is sufficient to preserve guarantees. It is also strongly influenced by the discount factor $\gamma$, which defines an implicit horizon. Smaller values permit larger deviations from $\pi_{\text{b}}$ and relax the accuracy required of the world model.
Safe policy improvement. We consider the setting where the world model is used to improve the behavioral policy $\pi_{\text{b}}=\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}\circ\phi$, with $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}\in\mkern 1.5mu\overline{\mkern-1.5mu\Pi\mkern-1.5mu}\mkern 1.5mu$ and the representation $\phi$ is fixed during each update. Restricting updates to a well-defined neighborhood guarantees that $\rho\mathopen{}\mathclose{{\left(\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\circ\phi,\mathcal{M}}}\right)-\rho\mathopen{}\mathclose{{\left(\pi_{\text{b}},\mathcal{M}}}\right)\geq\rho\mathopen{}\mathclose{{\left(\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu,\,\overline{\!{\mathcal{M}}}}}\right)-\rho\mathopen{}\mathclose{{\left(\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b},\,\overline{\!{\mathcal{M}}}}}\right)-\zeta$, where $\zeta$ is defined as the cumulative modeling error from the local losses.
Theorem 3.
(Deep, Safe Policy Improvement) Under the same preamble as in Thm. 2, assume that $\phi$ if fixed during the policy update and the behavioral is a latent policy with $\pi_{\text{b}}\coloneqq\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}\circ\phi$ and $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}\in\mkern 1.5mu\overline{\mkern-1.5mu\Pi\mkern-1.5mu}\mkern 1.5mu$. Then, the improvement of the return of $\mathcal{M}$ under $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$ can be guaranteed on $\pi_{\text{b}}$ as
$\rho\mathopen{}\mathclose{{\left(\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\circ\phi,\mathcal{M}}}\right)-\rho\mathopen{}\mathclose{{\left(\pi_{\text{b}},\mathcal{M}}}\right)\geq\rho\mathopen{}\mathclose{{\left(\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu,\,\overline{\!{\mathcal{M}}}}}\right)-\rho\mathopen{}\mathclose{{\left(\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b},\,\overline{\!{\mathcal{M}}}}}\right)-\zeta,\\ \text{ where }\zeta\coloneqq\textsc{AEL}\mathopen{}\mathclose{{\left(\pi_{\text{b}}}}\right)\cdot\mathopen{}\mathclose{{\left(\nicefrac{{L_{{R}}^{\xi_{\pi_{\text{b}}}}}}{{\gamma}}+K_{V}L_{P}^{\xi_{\pi_{\text{b}}}}}}\right)\mathopen{}\mathclose{{\left(\frac{1}{\nicefrac{{1}}{{D^{\sup}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi_{\text{b}},\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}}\right)}}-\gamma}+\frac{1}{1-\gamma}}}\right).$
Theorem 3 addresses the OOT issue (Section 3.1): if the SIR of the behavioral remains strictly below $\nicefrac{{1}}{{\gamma}}$, then minimizing the local losses reduces the error $\zeta$, ensuring safe policy improvement when the world model is used to enhance the policy. While our focus is not on offline SPI, Appendix E (Thm. LABEL:thm:deep-spi-pac) additionally provides a PAC variant of the result, following the standard use of confidence bounds in the SPI literature.
Representation learning. Finally, we analyze how learning a world model using our loss functions as an auxiliary task facilitates the learning of a useful representation. A good representation should ensure that environment states that are close in the representation also have close values, directly supporting policy learning. Specifically, we seek “almost” Lipschitz continuity (Vanderbei, 1991) of the form $\exists K\colon\forall s_{1},s_{2}\in\mathcal{S},\left|V^{\pi_{\text{b}}}\mathopen{}\mathclose{{\left(s_{1}}}\right)-V^{\pi_{\text{b}}}\mathopen{}\mathclose{{\left(s_{2}}}\right)\right|\leq K\cdot\,\overline{\!{d}}\mathopen{}\mathclose{{\left(\phi_{\text{old}}\mathopen{}\mathclose{{\left(s_{1}}}\right),\phi_{\text{old}}\mathopen{}\mathclose{{\left(s_{2}}}\right)}}\right)+\mathcal{L}_{\pi_{\text{b}}}\mathopen{}\mathclose{{\left(\phi_{\text{old}}}}\right)$ where $\mathcal{L}_{\pi_{\text{b}}}$ is an auxiliary loss depending on the data collected by $\bm{\pi_{\text{b}}}$. Notably, a critical question is whether updating the policy and its representation, respectively to $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$ and $\phi$, maintains Lipschitz continuity. Crucially, as the behavioral $\pi_{\text{b}}$ is updated to $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\circ\phi$ with respect to the experience collected under $\pi_{\text{b}}$, the bound must hold for $\mathcal{L}_{\pi_{\text{b}}}$. The following theorem is a probabilistic version of this statement, formalized as a concentration inequality:
Theorem 4.
(Deep SPI for representation learning) Under the same preamble as in Thm. 2, let $\varepsilon>0$ and $\delta\coloneqq 4\cdot\frac{L_{{R}}^{\xi_{\pi_{\text{b}}}}+\gamma K_{V}\cdot L_{P}^{\xi_{\pi_{\text{b}}}}}{\varepsilon\cdot\mathopen{}\mathclose{{\left(\nicefrac{{1}}{{D^{\sup}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi_{\text{b}},\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}}\right)}}-\gamma}}\right)}.$ Then, with probability at least $1-\delta$ under $\xi_{\pi_{\text{b}}}$, we have for all $s_{1},s_{2}\in\mathcal{S}$ that
$\left|V^{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}\mathopen{}\mathclose{{\left(s_{1}}}\right)-V^{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}\mathopen{}\mathclose{{\left(s_{2}}}\right)\right|\leq K_{V}\cdot\,\overline{\!{d}}\mathopen{}\mathclose{{\left(\phi\mathopen{}\mathclose{{\left(s_{1}}}\right),\phi\mathopen{}\mathclose{{\left(s_{2}}}\right)}}\right)+\varepsilon.$
Theorem 4 addresses confounding policy updates (Section 3.2): minimizing the losses increases the probability that learned representations remain almost Lipschitz under controlled policy changes (with an SIR below $\nicefrac{{1}}{{\gamma}}$). This prevents distinct states from collapsing into identical latent representations that degrade performance. We note that Gelada et al. (2019) proved a similar bound when $\pi_{\text{b}}=\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$ (the policy update was disregarded), which in contrast to ours, surely holds with
$\varepsilon\coloneqq\frac{L_{{R}}^{\xi_{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}}+\gamma K_{V}\cdot L_{P}^{\xi_{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}}}{1-\gamma}\cdot\mathopen{}\mathclose{{\left(\frac{1}{\xi_{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}\mathopen{}\mathclose{{\left(s_{1}}}\right)}+\frac{1}{\xi_{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}\mathopen{}\mathclose{{\left(s_{2}}}\right)}}}\right).$
However, in general spaces, for any specific $s\in\mathcal{S}$, $\xi_{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}\mathopen{}\mathclose{{\left(s}}\right)$ might simply equal zero, making the bound undefined. In particular, in the continuous setting, $\mathcal{S}$ is widely assumed to be endowed with a Borel sigma-algebra, where the probability of every single point is indeed zero.
6 Across the SPI-verse: PPO comes into play
These theorems inspire a practical RL algorithm that combines policy improvement and guarantees with solid empirical performance. The critical part of our approach is to make sure updates are restricted to the policy neighborhood while minimizing the auxiliary losses $L_{{R}}$, $L_{P}$. In fact, our neighborhood operator has close connections to PPO (Schulman et al., 2017), where the policy update is given by333we give the formulation of Kuba et al. (2022), which is equal to the one of Schulman et al. (2017).
$\pi_{n+1}$ $\coloneqq\operatorname*{arg\,sup}_{\pi^{\prime}\in\Pi}\mathop{\mathbb{E}}_{s\sim\xi_{\pi_{n}}}\left[\mathop{\mathbb{E}}_{a\sim\pi^{\prime}\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)}{A^{\pi_{n}}\mathopen{}\mathclose{{\left(s,a}}\right)}-\mathfrak{D}_{\pi_{n}}\mathopen{}\mathclose{{\left(\pi^{\prime}\mid s}}\right)\right],$ (5)
with $\mathfrak{D}_{\pi_{n}}\mathopen{}\mathclose{{\left(\pi^{\prime}\mid s}}\right)=\mathop{\mathbb{E}}_{a\sim\pi_{n}\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)}{\text{ReLu}\Big({\big[{\nicefrac{{\pi^{\prime}\mathopen{}\mathclose{{\left(a\mid s}}\right)}}{{\pi_{n}\mathopen{}\mathclose{{\left(a\mid s}}\right)}}-\text{clip}\mathopen{}\mathclose{{\left(\nicefrac{{\pi^{\prime}\mathopen{}\mathclose{{\left(a\mid s}}\right)}}{{\pi_{n}\mathopen{}\mathclose{{\left(a\mid s}}\right)}},\,1\pm\epsilon}}\right)}\big]\cdot A^{\pi_{n}}\mathopen{}\mathclose{{\left(s,a}}\right)}}\Big)$, for some $\epsilon>0$. By fixing $\epsilon=C-1$, instead of strictly constraining the updates to the neighborhood, the regularization $\mathfrak{D}_{\pi_{n}}\mathopen{}\mathclose{{\left(\pi^{\prime}\mid s}}\right)$ corrects the utility $\mathop{\mathbb{E}}_{a\sim\pi^{\prime}\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)}A^{\pi_{n}}\mathopen{}\mathclose{{\left(s,a}}\right)$ (compare Eq. 3 and Eq. 5), so that there is no incentive for $\pi^{\prime}$ to deviate from $\pi_{n}$ with an IR outside the range $\mathopen{[}2-C,C\mathclose{]}$. Under the same assumption as in Theorem 1, PPO is also an instance of mirror learning (Kuba et al., 2022), meaning it also benefits from the same convergence guarantees.
Strictly restricting the IR in a neighborhood is much harder in practice, considering a PPO objective is thus an appealing alternative. However, it is not sufficient to add the auxiliary losses $L_{P},L_{R}$ to the objective of Eq. 5 to maintain the guarantees. Indeed, updating the representation $\phi$ by minimizing the additional losses may push the the policy $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\circ\phi$ outside the neighborhood. As a solution we propose to incorporate the local losses by replacing all occurrences of $A^{\pi_{n}}$ in Eq. 5 by the utility
$U^{\pi_{n}}\mathopen{}\mathclose{{\left(s,a,s^{\prime}}}\right)\coloneqq A^{\pi_{n}}\mathopen{}\mathclose{{\left(s,a}}\right)-\alpha_{R}\cdot\ell_{R}\mathopen{}\mathclose{{\left(s,a}}\right)-\alpha_{P}\cdot\ell_{P}\mathopen{}\mathclose{{\left(s,a,s^{\prime}}}\right),$ (6)
where $\ell_{R}\mathopen{}\mathclose{{\left(s,a}}\right)\coloneqq\left|{R}\mathopen{}\mathclose{{\left(s,a}}\right)-\,\overline{\!{{R}}}\mathopen{}\mathclose{{\left(\phi\mathopen{}\mathclose{{\left(s}}\right),a}}\right)\right|$, $\ell_{P}\mathopen{}\mathclose{{\left(s,a,s^{\prime}}}\right)\coloneqq\mathop{\mathbb{E}}_{\bar{s}^{\prime}\sim\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left({\cdotp}\mid\phi\mathopen{}\mathclose{{\left(s}}\right),a}}\right)}\,\overline{\!{d}}\mathopen{}\mathclose{{\left(\phi\mathopen{}\mathclose{{\left(s^{\prime}}}\right),\bar{s}^{\prime}}}\right)$, $s^{\prime}\sim P\mathopen{}\mathclose{{\left({\cdotp}\mid s,a}}\right)$, and $\alpha_{R},\alpha_{P}\in\mathopen{(}0,1\mathclose{]}$. Intuitively, $\ell_{R}$, $\ell_{P}$ are transition-wise auxiliary losses that allow retrieving $L_{{R}}^{\xi_{\pi_{n}}}$ and $L_{P}^{\xi_{\pi_{n}}}$ in expectation w.r.t. the current policy $\pi_{n}$. When optimized, since they are clipped in a PPO-fashion, $U^{\pi_{n}}$ allows restricting the policy updates to the neighborhood.
Inputs: Horizon $T$, batch size $B$, vectorized environment env, parameters $\theta$
Initialize vectors $\bm{s}\in\mathcal{S}^{\mathopen{}\mathclose{{\left(T+1}}\right)\times B},\bm{a}\in\mathcal{A}^{T\times B},\bm{r}\in\mathbb{R}^{T\times B}$
repeat
 for $t\leftarrow 1$ to $T$ do
      Draw actions from the current policy: $\bm{a}_{t,i}\sim\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left({\cdotp}\mid\phi\mathopen{}\mathclose{{\left(\bm{s}_{t,i}}}\right)}}\right)$ $\forall 1\leq i\leq B$
      Perform a single parallelized ($B$) step: $\bm{r}_{t},\bm{s}_{t+1}\leftarrow\texttt{env.step}\mathopen{}\mathclose{{\left(\bm{s}_{t},\bm{a}_{t}}}\right)$
   end for
  Update $\theta$ by descending $\nabla_{\theta}\,{\texttt{DeepSPI\_loss}}(\bm{s},\bm{a},\bm{r},U^{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\circ\phi},\,\theta)$
 $\triangleright$ change $A$ in Eq. 5 by $U$ from Eq. 6
 $\bm{s}_{1}\leftarrow\bm{s}_{T+1}$
until convergence
return $\theta$
Algorithm 1 DeepSPI
From this loss, we propose DeepSPI, a principled algorithm leveraging the policy improvement and representation learning capabilities developed in our theory. As our losses rely on distributions defined over the current policy, we focus on the on-policy setting. While model-based approaches are not standard in this setting, we stress that highly parallelized collection of data (e.g., via vectorized environments) enables a wide coverage of the state space (cf. Mayor et al., 2025; Gallici et al., 2025), which is suitable to optimize the latent model. DeepSPI updates the world model, the encoder, and the policy simultaneously while guaranteeing the representation is suited to perform safe policy updates.
6.1 Illustrative Example
Figure 3: Toy maze environment illustrating the confounding policy update problem.
To illustrate the representation learning capabilities of DeepSPI, we consider the toy grid-world shown in Fig. 3. This environment mirrors the confounding policy update discussed in Sect. 3.2, instantiated earlier in Fig. 2.
The agent starts in the cell labeled I. Upon leaving the orange cell immediately to its right, it is sent to the top branch with probability $1-\epsilon$ and to the bottom branch with probability $\epsilon$. It must then traverse a corridor of $n$ blue cells (here nn=5). Moving one cell to the right yields a reward of +1, and the agent cannot move backwards.
At the final corridor cell, marked with a $\star$, moving right yields a reward of +1 regardless of whether the agent is in the top or bottom branch, and the episode terminates. The difference is when the agent moves up from the $\star$ cell: in the top branch, it receives a reward of $\nicefrac{{+n}}{{\gamma^{n}}}$, whereas in the bottom branch it receives $\nicefrac{{-(2-\epsilon)n}}{{\mathopen{}\mathclose{{\left(\epsilon\,\gamma^{n}}}\right)}}$ before termination. As in Sect. 3.2, this construction ensures that if both $\star$ states are merged in the latent space, choosing “right” remains acceptable (their values coincide), but choosing “up” produces a negative expected return from the initial state I (details in Appendix LABEL:appendix:toy-example). To improve upon the policy that always chooses “right,” the agent must learn to assign distinct representations to the two $\star$ cells, select “up” in the top one, and “right” in the bottom one.
We compare the behaviour of PPO and DeepSPI in this environment. Since our goal is to highlight the agent’s representation-learning capabilities, each observation is provided as raw pixels.
Figure 4: Value from cell I in the maze (left) and distance between the representation of the $\star$ cell from the top and bottom branches (right).
The agent must therefore learn both a policy and an encoder mapping pixels to a structured latent space. Further details on the environment and observation scheme are given in Appendix LABEL:appendix:toy-example. As shown in Fig. 4, the representation learned by PPO collapses the top and bottom $\star$ cells into a single latent state. With such a representation, the best policy PPO can learn is to always choose “right,” which leads to a return of $\sim\!4.8$. In contrast, DeepSPI benefits from the representation quality guarantees of Thm. 4, which ensure that states with different values remain separated in the latent space for all policies in a suitable neighborhood. This is exactly what we observe: the learned representation distinguishes the two $\star$ cells. As a result, the agent learns to choose “up” in the top $\star$ cell and “right” in the bottom one, achieving a return of $\sim\!8$.
7 Experiments
Figure 5: Aggregate results on stochastic versions of the standard 57 environments from ALE, with 95% confidence intervals (CIs). Higher values for the mean, median, and interquartile mean (IQM) indicate better performance, while a lower optimality gap is preferable (cf. Agarwal et al. 2021b). CIs are obtained through percentile bootstrapping with stratified resampling. Plots per environment available in Appendix LABEL:appendix:full-plots.
Figure 6: Sample efficiency w.r.t. IQM normalized scores on the stochastic ALE-57. Shaded regions give pointwise 95% CIs obtained via percentile stratified bootstrap.
In this section, we evaluate the practical performance of DeepSPI in environments where (i) representation learning is essential and (ii) dynamics are complex. We use the Atari Arcade Learning Environment (ALE; Bellemare et al. 2013) and consider each state as four stacked frames. ALE domains feature a wide range of dynamics; to further introduce stochasticity, we follow Machado et al. (2018) and employ two standard tricks: sticky actions, where with probability $p_{a}$ the previous action is repeated (simulating joystick or reaction-time imperfections), and random initialisation, where the agent begins after $n_{\text{NOOP}}$ initial no-op frames. We set $p_{a}=0.3$ and $n_{\text{NOOP}}=60$.
As baselines, we consider PPO (vectorized cleanRL implementation; Huang et al., 2022) and DeepMDPs (Gelada et al., 2019). Essentially, DeepMDPs are principled auxiliary tasks (the losses $L_{{R}},L_{P}$ presented in Sect. 5) that can be plugged to any RL algorithm to improve the representation learned (with guarantees). The main difference with DeepSPI is that $L_{{R}}$, $L_{P}$ are able to push the updated policy out of the neighborhood by learning the representation via the additional losses, for which updates are not constrained. This means that none of the SPI guarantees presented in this paper apply to DeepMDPs. For a fair comparison, we plugged the DeepMDP losses to (vectorized) PPO, and we use the architecture as for DeepSPI. We use the default cleanRL’s hyperparameters for the three algorithms, except for the data collection ($128$ environments with a horizon of $8$ steps).
As latent space, we use the raw 33D representation obtained after the convolution layers (as recommended and used by Gelada et al., 2019). For modeling the transition function, we found best to use a mixture of multivariate normal distributions (the transition network outputs $5$ means/diagonal matrices). To deal with the Lipschitz constraints that need to be enforced on the reward and transition functions, we found the most efficient to model $\overline{\!{{R}}},\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu$ via Lipschitz networks (precisely, we use norm-constrained GroupSort architectures to enforce $1$-Lipschitzness; Anil et al., 2019).
As shown in Fig. 5 and 6, DeepSPI delivers strong performance, improving on both PPO and DeepMDP. Notably, these results are obtained while preserving SPI-style properties; a valuable combination, as such theoretical control typically comes at the expense of performance and substantial data requirements. Beyond pure performance, we want to assess whether the world model, learned via DeepSPI, exhibits accurate dynamics.
Figure 7: Median transition and reward losses during training, aggregated across all the ALE. For the sake of visualization, we cut $L_{P}$ lower values from the plot.
Fig. 7 reports $L_{P}$, $L_{{R}}$ during training. Note that DeepSPI consistently achieves lower transition losses, indicating more accurate transition functions. We discuss the statistical significance of that statement in Appendix LABEL:appendix:full-plots. In contrast to the off-policy setting where Gelada et al. (2019) reported competing transition and reward losses, we did not observe such behavior in our parallel on-policy setting. We attribute this stability to the fact that our losses are always computed under the current policy, unlike off-policy methods that rely on replay buffers.
To probe the predictive quality of the latent model and illustrate Thm. 3, we introduced DreamSPI, a naïve variant where DeepSPI learns the world model and representation, and PPO updates the policy from imagined trajectories (Appendix LABEL:appendix:dream-spi). Unlike off-policy approaches that exploit replay buffers and update the model at every interaction step, our fully on-policy setting updates the world model only from fresh interaction data, which makes combining model learning and planning inherently more challenging. Even so, DreamSPI achieves learning progress in several environments and exhibits coherent behaviours (cf. Fig. 8 & Appendix LABEL:appendix:full-plots). While its aggregate median score remains below the baselines, this is somehow expected given the stricter data requirements compared to usual model-based approaches. Importantly, the ability to maintain a model offers benefits that extend far beyond raw scores, enabling future applications in safety, verification, and reactive synthesis.
Figure 8: Sample environments from ALE where DreamSPI learns meaningful behaviors.
8 Conclusion and Future Work
We developed a theoretical framework for safe policy improvement (SPI) that combines world-model and representation learning in nontrivial settings. Our results show that constraining policy updates within a well-defined neighborhood yields monotonic improvement and convergence, while auxiliary transition and reward losses ensure that the latent space remains suitable for policy optimisation. We further provided model-quality guarantees in the form of a “deep” SPI theorem, which jointly accounts for the learned representation and the reward/transition losses. These results directly address two critical issues in model-based RL: out-of-trajectory errors and confounding policy updates. Building on this analysis, we proposed DeepSPI, a principled algorithm that integrates the theoretical ingredients with PPO. On ALE, DeepSPI is competitive with and often improves upon PPO and DeepMDPs, while providing SPI guarantees.
This work opens several directions. A first avenue is to make pure deep SPI model-based planning practical. Our experiments with DreamSPI suggest that this is feasible but requires improved sample efficiency. Another direction goes beyond return optimization: a principled world model, grounded in our theory, can support safe reinforcement learning via formal methods, through synthesis (Delgrange et al., 2025; Lechner et al., 2022), or shielding (Jansen et al., 2020).
Acknowledgments
We thank Marnix Suilen and Guillermo A. Pérez for their valuable feedback during the preparation of this manuscript. This research was supported by the Belgian Flemish AI Research Program and the “DESCARTES” iBOF project. W. Röpke and R. Avalos are supported by the Research Foundation – Flanders (FWO), with respective grant numbers 1197622N and 11F5721N.
Appendix
Appendix A Remark on value functions and episodic processes
An episodic process is formally defined as an MDP $\mathcal{M}=\langle\mathcal{S},\mathcal{A},P,{R},s_{\mathit{I}},\gamma\rangle$ where:
  1. (i)
    there is a special state $s_{\textit{reset}}\in\mathcal{S}$, intuitively indicating the termination of any episode;
  2. (ii)
    the reset state does not incur any reward: ${R}\mathopen{}\mathclose{{\left(s_{\textit{reset}},a}}\right)=0$ for all actions $a\in\mathcal{A}$;
  3. (iii)
    $s_{\textit{reset}}$ is almost surely visited under any policy: for all policies $\pi\in\Pi$, $\mathbb{P}_{\pi}\mathopen{}\mathclose{{\left(\left\{\mathopen{}\mathclose{{\left(s_{t},a_{t}}}\right)_{t\geq 0}\mid\exists i\colon s_{i}=s_{\textit{reset}}\right\}}}\right)=1$; and
  4. (iv)
    $\mathcal{M}$ restarts from the initial state once reset: $P\mathopen{}\mathclose{{\left(\left\{s_{\mathit{I}}\right\}\mid s_{\textit{reset}},a}}\right)=1$ for all $a\in\mathcal{A}$.
Note that by items iii and iv, $s_{\textit{reset}}$ is almost surely infinitely often visited: we have for all $\pi\in\Pi$ that
$\mathbb{P}_{\pi}\mathopen{}\mathclose{{\left(\left\{\mathopen{}\mathclose{{\left(s_{t},a_{t}}}\right)_{t\geq 0}\mid\forall i\geq 0,\,\exists j>i\colon s_{j}=s_{\textit{reset}}\right\}}}\right)=1.$
Alternatively and equivalently, an episodic process may also be defined without a unique reset state by the means of several terminal states, which go back to the initial state with probability one.
An episode of $\mathcal{M}$ is thus the prefix $s_{0},a_{0},\dots,a_{t-1},s_{t}$ of a trajectory where $s_{t}=s_{\textit{reset}}$ and for all $i<t$, $s_{i}\neq s_{\textit{reset}}$. Notice that our formulation embeds (but is not limited to) finite-horizon tasks, where an upper bound on the length of the episodes is fixed. The average episode length (AEL) of $\pi$ is then formally defined as $\textsc{AEL}\mathopen{}\mathclose{{\left(\pi}}\right)=\displaystyle\mathbb{E}_{\pi}\left[\mathbf{T}\right]$ with
$\mathbf{T}\mathopen{}\mathclose{{\left(\tau}}\right)=\sum_{i=0}^{\infty}\mathopen{}\mathclose{{\left(i+1}}\right)\cdot\mathds{1}\left\{s_{i}=s_{\textit{reset}}\,\text{ and }\,\forall j<i,\,s_{j}\neq s_{\textit{reset}}\right\}$
for any trajectory $\tau=\mathopen{}\mathclose{{\left(s_{t},a_{t}}}\right)_{t\geq 0}$.
Often, when considering episodic tasks, RL algorithms stops accumulating rewards upon the termination of every episode. In practical implementations, this corresponds to discarding rewards when a flag done, indicating episode termination, is set to true. In such case, we may slightly adapt our value functions as:
$V^{\pi}\mathopen{}\mathclose{{\left(s}}\right)$ $=\begin{cases}\displaystyle\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\mathopen{}\mathclose{{\left(\prod_{i=1}^{t}\mathds{1}\left\{s_{i}\neq s_{\textit{reset}}\right\}\cdot\gamma}}\right){R}\mathopen{}\mathclose{{\left(s_{t},a_{t}}}\right)\,\Big|\,s_{0}=s\right]&\text{if }s\neq s_{\textit{reset}}\\ 0&\text{otherwise};\end{cases}$
or, when formalized as Bellman’s equation:
$Q^{\pi}\mathopen{}\mathclose{{\left(s,a}}\right)$ $=\begin{cases}{R}\mathopen{}\mathclose{{\left(s,a}}\right)+\gamma\cdot\mathop{\mathbb{E}}_{s^{\prime}\sim P\mathopen{}\mathclose{{\left({\cdotp}\mid s,a}}\right)}V^{\pi}\mathopen{}\mathclose{{\left(s^{\prime}}}\right)&\text{if }s\neq s_{\textit{reset}}\\ 0&\text{otherwise; and}\end{cases}$
$V^{\pi}\mathopen{}\mathclose{{\left(s}}\right)$ $=\mathop{\mathbb{E}}_{a\sim\pi\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)}Q^{\pi}\mathopen{}\mathclose{{\left(s,a}}\right).$
All our results extend to this formulation (cf. Remark LABEL:rmk:extension-lem-episodic).
Remark 1 (Occupancy measure).
In RL theory, the discounted occupancy measure
$\mu^{\gamma}_{\pi}\mathopen{}\mathclose{{\left(s}}\right)\coloneqq\mathopen{}\mathclose{{\left(1-\gamma}}\right)\cdot\sum_{t=0}^{\infty}\gamma^{t}\mathbb{P}_{\pi}\mathopen{}\mathclose{{\left(\left\{\mathopen{}\mathclose{{\left(s_{i},a_{i}}}\right)_{i\geq 0}\mid s_{t}=s\right\}}}\right)$
is often considered as the default marginal distribution over states the agent visit along the interaction, mostly because of its suitable theoretical properties. In fact, for any arbitrary MDP, $\mu^{\gamma}_{\pi}$ is the stationary distribution of the episodic process obtained by considering a reset probability of $1-\gamma$ from every state of the original MDP (Puterman, 1994; Metelli et al., 2023). Again, we contend that all our results can be extended to the occupancy measure with little effort.
Appendix B Policy Improvements through Mirror Learning and Convergence Guarantees
In this section, we prove that $\mathcal{N}^{C}$ (Eq. 2) is a proper mirror learning neighborhood operator. As a consequence, appropriately updating the policy according to $\mathcal{N}^{C}$ is guaranteed to be an instance of mirror learning, yielding the convergence guarantees of Theorem 1.
For completeness, we recall the definition of neighborhood operator from Kuba et al. (2022).
Definition 1 (Neighborhood operator).
The mapping $\mathcal{N}\colon\Pi\to 2^{\Pi}$ is a (mirror learning) neighborhood operator, if
  1. 1.
    (continuity) It is a continuous map;
  2. 2.
    (compactness) Every $\mathcal{N}\mathopen{}\mathclose{{\left(\pi}}\right)$ is a compact set; and
  3. 3.
    (closed ball) There exists a metric $d\colon\Pi\times\Pi\to\mathopen{[}0,\infty\mathclose{)}$, such that for all policies $\pi\in\Pi$, there exists $\epsilon>0$, such that $d\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)\leq\epsilon$ implies $\pi^{\prime}\in\mathcal{N}\mathopen{}\mathclose{{\left(\pi}}\right)$.
The trivial neighborhood operator is $\mathcal{N}\mathopen{}\mathclose{{\left(\pi}}\right)=\Pi$.
Lemma 1.
$\mathcal{N}^{C}$ is a neighborhood operator.
Proof.
Henceforth, fix a policy $\pi\in\Pi$. When taking the supremum, infimum, maximum, or minimum value over states and actions, we always consider actions to be taken from the support of the behavioral policy (in the denominator of the quotient).
Item 2 (compactness) is trivial due to $D^{\inf}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)\geq 2-C$ and $D^{\sup}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)\leq C$ for any $\pi^{\prime}\in\mathcal{N}^{C}\mathopen{}\mathclose{{\left(\pi}}\right)$. This means $\mathcal{N}^{C}\mathopen{}\mathclose{{\left(\pi}}\right)$ contains its extrema, i.e., all the policies $\pi^{\prime}$ satisfying $D^{\inf}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)=2-C$ and $D^{\sup}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)\leq C$, or $D^{\inf}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)\geq 2-C$ and $D^{\sup}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)=C$.
In the following, for any $\pi\in\Pi$ and sequence $\mathopen{}\mathclose{{\left(\pi_{n}}}\right)_{n\geq 0}$, we write $\pi_{n}\to\pi$ for the convergence of the sequence to $\pi$ with respect to the metric
$d\mathopen{}\mathclose{{\left(\pi_{1},\pi_{2}}}\right)=\begin{cases}\left\|\pi_{1}-\pi_{2}\right\|_{\infty}&\text{if }\text{supp}\mathopen{}\mathclose{{\left(\pi_{1}\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)}}\right)=\text{supp}\mathopen{}\mathclose{{\left(\pi_{2}\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)}}\right)\quad\forall s\in\mathcal{S},\text{ and}\\ 1&\text{otherwise}.\end{cases}$ (7)
In other words, $\pi_{n}\to\pi$ means that $\pi_{n}$ converges to $\pi$ in supremum norm as $n\to\infty$ when the support of the converging policy stabilizes and becomes the same as the limit policy.
Let us prove item 1 (continuity). We show that $\mathcal{N}^{C}$ is a continuous correspondence by showing it is upper and lower hemicontinuous (Ok, 2007).
$\mathcal{N}^{C}$ is upper hemicontinuous (uhc) if it is compact-valued (item 1) and, for all policies $\pi\in\Pi$ and every sequences $\mathopen{}\mathclose{{\left(\pi_{n}}}\right)_{n\geq 0}$ and $\mathopen{}\mathclose{{\left(\pi^{\prime}_{n}}}\right)_{n\geq 0}$ with $\pi^{\prime}_{n}\in\mathcal{N}^{C}\mathopen{}\mathclose{{\left(\pi_{n}}}\right)$ for all $n\geq 0$, $\pi_{n}\to\pi$ and $\pi^{\prime}_{n}\to\pi^{\prime}$ implies $\pi^{\prime}\in\mathcal{N}^{C}\mathopen{}\mathclose{{\left(\pi}}\right)$. Let $\mathopen{}\mathclose{{\left(\pi_{n}}}\right)_{n\geq 0}$ and $\mathopen{}\mathclose{{\left(\pi^{\prime}_{n}}}\right)_{n\geq 0}$ be sequences of policies with $\pi^{\prime}_{n}\in\mathcal{N}^{C}\mathopen{}\mathclose{{\left(\pi_{n}}}\right)$ for all $n\geq 0$.
Fix $s\in\mathcal{S}$ and $a\in\mathcal{A}$. Consider the mapping
$f_{s,a}\colon\left\{\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)\in\Pi\times\Pi\mid a\in\text{supp}\mathopen{}\mathclose{{\left(\pi\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)}}\right)\right\}\to\mathopen{[}0,\infty),\quad\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)\mapsto\frac{\pi^{\prime}\mathopen{}\mathclose{{\left(a\mid s}}\right)}{\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)}.$
It is clear $f_{s,a}$ is continuous since the application of $\pi$ to $\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)$ is continuous and the division of two continuous functions is also continuous (when considering actions from the support of $\pi\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)$). Importantly, for $\operatorname*{ext}\in\left\{\sup,\inf\right\}$, $D^{\operatorname*{ext}}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)=\operatorname*{ext}\left\{f_{s,a}\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)\colon s\in\mathcal{S},a\in\text{supp}\mathopen{}\mathclose{{\left(\pi\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)}}\right)\right\}$ is also continuous: since $\mathcal{S}$ and $\mathcal{A}$ are finite, the supremum (resp. infimum) boils down to taking the maximum (resp. minimum) of finitely many many continuous functions, which is a continuous operation.
Now, assume that $\pi_{n}\to\pi$ and $\pi^{\prime}_{n}\to\pi^{\prime}$. The continuity of $D^{\operatorname*{ext}}_{\text{IR}}$ means that $D^{\operatorname*{ext}}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi_{n},\pi^{\prime}_{n}}}\right)\to D^{\operatorname*{ext}}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)$. Since $\pi^{\prime}_{n}\in\mathcal{N}^{C}\mathopen{}\mathclose{{\left(\pi_{n}}}\right)$, we have $D^{\inf}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi_{n},\pi^{\prime}_{n}}}\right)\geq 2-C$ and $D^{\sup}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi_{n},\pi^{\prime}_{n}}}\right)\leq C$ for all $n\geq 0$. By the fact that $D^{\operatorname*{ext}}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi_{n},\pi^{\prime}_{n}}}\right)$ converges to $D^{\operatorname*{ext}}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)$ for $\operatorname*{ext}\in\left\{\inf,\sup\right\}$, we also have that $D^{\inf}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)\geq 2-C$ and $D^{\sup}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)\leq C$.
Then, $\mathcal{N}^{C}$ is uhc.
$\mathcal{N}^{C}$ is lower hemicontinuous (lhc) if, for every policy $\pi$, sequence $\mathopen{}\mathclose{{\left(\pi_{n}}}\right)_{n\geq 0}$ with $\pi_{n}\to\pi$, and policy $\pi^{\prime}\in\mathcal{N}^{C}\mathopen{}\mathclose{{\left(\pi}}\right)$, there exists a sequence $\mathopen{}\mathclose{{\left(\pi^{\prime}_{n}}}\right)_{n\geq 0}$ with $\pi^{\prime}_{n}\to\pi^{\prime}$ and such that there is a $n_{0}\geq 0$ from which, for all $n\geq n_{0}$, $\pi^{\prime}_{n}\in\mathcal{N}^{C}\mathopen{}\mathclose{{\left(\pi_{n}}}\right)$.
Therefore, let $\mathopen{}\mathclose{{\left(\pi_{n}}}\right)_{n\geq 0}$ be a sequence of policies so that $\pi_{n}\to\pi$ and $\pi^{\prime}\in\mathcal{N}\mathopen{}\mathclose{{\left(\pi}}\right)$. Since $\pi_{n}\to\pi$, we have
$\forall\delta>0,\exists n_{0}\in\mathbb{N}\colon\forall n\geq n_{0},\left\|\pi_{n}-\pi\right\|_{\infty}\leq\delta\;\text{ and }\;\text{supp}\mathopen{}\mathclose{{\left(\pi_{n}\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)}}\right)=\text{supp}\mathopen{}\mathclose{{\left(\pi\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)}}\right)\quad\forall s\in\mathcal{S}.$
In particular, this holds for $\delta<\nicefrac{{\pi_{\min}}}{{2}}$, where $\pi_{\min}=\min\left\{\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)\colon s\in\mathcal{S},a\in\text{supp}\mathopen{}\mathclose{{\left(\pi\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)}}\right)\right\}.$ Let $n_{0}\geq 0$ be the step associated with $\delta<\nicefrac{{\pi_{\min}}}{{2}}$ and $n\geq n_{0}$. Write $\delta_{n}=\left\|\pi_{n}-\pi\right\|_{\infty}$ and let
$\epsilon_{n}=\frac{2C\delta_{n}}{\pi_{\min}\mathopen{}\mathclose{{\left(C-1}}\right)+2C\delta_{n}}\in\mathopen{(}0,1\mathclose{)}$
Construct a sequence $\mathopen{}\mathclose{{\left(\pi^{\prime}_{n}}}\right)_{n\geq 0}$ so that, for all $s\in\mathcal{S}$, $a\in\mathcal{A}$, and $n\geq n_{0}$,
$\pi^{\prime}_{n}\mathopen{}\mathclose{{\left(a\mid s}}\right)=\mathopen{}\mathclose{{\left(1-\epsilon_{n}}}\right)\cdot\pi^{\prime}\mathopen{}\mathclose{{\left(a\mid s}}\right)+\epsilon_{n}\cdot\pi_{n}\mathopen{}\mathclose{{\left(a\mid s}}\right).$
Intuitively, $\pi^{\prime}_{n}$ is a mixture of distributions $\pi^{\prime}\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)$ and $\pi_{n}\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)$. Consequently, $\pi^{\prime}_{n}\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)$ is a well-defined distribution. Finally, note that $\pi^{\prime}_{n}\to\pi^{\prime}$ because $\delta_{n}\to 0$, and so does $\epsilon_{n}$.
Now, we restrict our attention to $a\in\text{supp}\mathopen{}\mathclose{{\left(\pi_{n}\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)}}\right)$. Note that since $\pi_{n}$ stably converges to $\pi$ with its support, $\pi$ has the same support as $\pi_{n}$. Furthermore, since $\pi^{\prime}\in\mathcal{N}^{C}\mathopen{}\mathclose{{\left(C}}\right)$, $\pi^{\prime}$ has also the same support as $\pi_{n}$. In consequence, $\pi^{\prime}_{n}$ has the same support as $\pi_{n}$.
Having that said, we start by showing the upper bound:
$\frac{\pi^{\prime}_{n}\mathopen{}\mathclose{{\left(a\mid s}}\right)}{\pi_{n}\mathopen{}\mathclose{{\left(a\mid s}}\right)}$ $=\mathopen{}\mathclose{{\left(1-\epsilon_{n}}}\right)\frac{\pi^{\prime}\mathopen{}\mathclose{{\left(a\mid s}}\right)}{\pi_{n}\mathopen{}\mathclose{{\left(a\mid s}}\right)}+\epsilon_{n}$
$\leq\mathopen{}\mathclose{{\left(1-\epsilon_{n}}}\right)\frac{C\cdot\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)}{\pi_{n}\mathopen{}\mathclose{{\left(a\mid s}}\right)}+\epsilon_{n}$ (because $\pi^{\prime}\mathopen{}\mathclose{{\left(a\mid s}}\right)\leq C\cdot\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)$)
$\leq\mathopen{}\mathclose{{\left(1-\epsilon_{n}}}\right)\cdot\frac{C\cdot\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)}{\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)-\delta_{n}}+\epsilon_{n}$ (because $\pi_{n}\mathopen{}\mathclose{{\left(a\mid s}}\right)\geq\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)-\delta_{n}$)
$=\mathopen{}\mathclose{{\left(1-\epsilon_{n}}}\right)\frac{C}{1-\nicefrac{{\delta_{n}}}{{\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)}}}+\epsilon_{n}$
$\leq\mathopen{}\mathclose{{\left(1-\epsilon_{n}}}\right)\frac{C}{1-\nicefrac{{\delta_{n}}}{{\pi_{\min}}}}+\epsilon_{n}.$
Note that for all $x\in\mathopen{[}0,\nicefrac{{1}}{{2}}\mathclose{]}$,
$\frac{1}{1-x}\leq 1+2x\;\text{ because }\;1+2x-\frac{1}{1-x}\geq 0\iff\frac{\mathopen{}\mathclose{{\left(1+2x}}\right)\mathopen{}\mathclose{{\left(1-x}}\right)-1}{1-x}\geq 0\iff\frac{x\mathopen{}\mathclose{{\left(1-2x}}\right)}{1-x}\geq 0.$
Then, since $0<\nicefrac{{\delta_{n}}}{{\pi_{\min}}}<\nicefrac{{1}}{{2}}$, we have
$\frac{\pi^{\prime}_{n}\mathopen{}\mathclose{{\left(a\mid s}}\right)}{\pi_{n}\mathopen{}\mathclose{{\left(\pi\mid s}}\right)}\leq\mathopen{}\mathclose{{\left(1-\epsilon_{n}}}\right)\cdot C\cdot\mathopen{}\mathclose{{\left(1+\nicefrac{{2\delta_{n}}}{{\pi_{\min}}}}}\right)+\epsilon_{n}.$
Let $x_{n}=1+\frac{2\delta_{n}}{\pi_{\min}}$, and note that
$\epsilon_{n}=\frac{2C\delta_{n}}{\pi_{\min}\mathopen{}\mathclose{{\left(C-1}}\right)+2C\delta_{n}}=\frac{2C\cdot\nicefrac{{\delta_{n}}}{{\pi_{\min}}}}{C+2C\cdot\nicefrac{{\delta_{n}}}{{\pi_{\min}}}-1}=\frac{-2C\cdot\nicefrac{{\delta_{n}}}{{\pi_{\min}}}}{1-C-2C\cdot\nicefrac{{\delta_{n}}}{{\pi_{\min}}}}=\frac{C\mathopen{}\mathclose{{\left(1-x_{n}}}\right)}{1-x_{n}\cdot C}.$
Then,
$\frac{\pi^{\prime}_{n}\mathopen{}\mathclose{{\left(a\mid s}}\right)}{\pi_{n}\mathopen{}\mathclose{{\left(a\mid s}}\right)}$ $\leq\mathopen{}\mathclose{{\left(1-\epsilon_{n}}}\right)x_{n}\cdot C+\epsilon_{n}$
$=x_{n}\cdot C-\epsilon_{n}\cdot x_{n}\cdot C+\epsilon_{n}$
$=x_{n}\cdot C-\frac{C\mathopen{}\mathclose{{\left(1-x_{n}}}\right)}{1-x_{n}\cdot C}\cdot x_{n}\cdot C+\frac{C\mathopen{}\mathclose{{\left(1-x_{n}}}\right)}{1-x_{n}\cdot C}$
$=\frac{x_{n}\cdot C\mathopen{}\mathclose{{\left(1-x_{n}\cdot C}}\right)-x_{n}\cdot C^{2}\mathopen{}\mathclose{{\left(1-x_{n}}}\right)+C\mathopen{}\mathclose{{\left(1-x_{n}}}\right)}{1-x_{n}\cdot C}$
$=\frac{x_{n}\cdot C-x_{n}^{2}C^{2}-x_{n}\cdot C^{2}+x_{n}^{2}C^{2}+C-x_{n}\cdot C}{1-x_{n}\cdot C}$
$=\frac{-x_{n}\cdot C^{2}+C}{1-x_{n}\cdot C}$
$=C\cdot\frac{1-x_{n}\cdot C}{1-x_{n}\cdot C}$
$=C,$
which means that $D^{\sup}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi_{n},\pi^{\prime}_{n}}}\right)\leq C$.
We now show the lower bound:
$\frac{\pi^{\prime}_{n}\mathopen{}\mathclose{{\left(a\mid s}}\right)}{\pi_{n}\mathopen{}\mathclose{{\left(a\mid s}}\right)}$ $=\mathopen{}\mathclose{{\left(1-\epsilon_{n}}}\right)\frac{\pi^{\prime}\mathopen{}\mathclose{{\left(a\mid s}}\right)}{\pi_{n}\mathopen{}\mathclose{{\left(a\mid s}}\right)}+\epsilon_{n}$
$\geq\mathopen{}\mathclose{{\left(1-\epsilon_{n}}}\right)\frac{\mathopen{}\mathclose{{\left(2-C}}\right)\cdot\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)}{\pi_{n}\mathopen{}\mathclose{{\left(a\mid s}}\right)}+\epsilon_{n}$ (because $\pi^{\prime}\mathopen{}\mathclose{{\left(a\mid s}}\right)\geq(2-C)\cdot\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)$)
$\geq\mathopen{}\mathclose{{\left(1-\epsilon_{n}}}\right)\frac{\mathopen{}\mathclose{{\left(2-C}}\right)\cdot\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)}{\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)+\delta_{n}}+\epsilon_{n}$ (because $\pi_{n}\mathopen{}\mathclose{{\left(a\mid s}}\right)\leq\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)+\delta_{n}$)
$=\mathopen{}\mathclose{{\left(1-\epsilon_{n}}}\right)\,\frac{\mathopen{}\mathclose{{\left(2-C}}\right)}{1+\nicefrac{{\delta_{n}}}{{\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)}}}+\epsilon_{n}$
$\geq\mathopen{}\mathclose{{\left(1-\epsilon_{n}}}\right)\,\frac{\mathopen{}\mathclose{{\left(2-C}}\right)}{1+\nicefrac{{\delta_{n}}}{{\pi_{\min}}}}+\epsilon_{n}$
$\geq\mathopen{}\mathclose{{\left(1-\epsilon_{n}}}\right)\cdot\mathopen{}\mathclose{{\left(2-C}}\right)\cdot\mathopen{}\mathclose{{\left(1-\nicefrac{{\delta_{n}}}{{\pi_{\min}}}}}\right)+\epsilon_{n}$ (because for all $x\in\mathbb{R},\,\frac{1}{1+x}\geq 1-x$)
$=\mathopen{}\mathclose{{\left(1-\epsilon_{n}}}\right)\cdot\mathopen{}\mathclose{{\left(2-C-2\cdot\nicefrac{{\delta_{n}}}{{{\pi_{\min}}}}+C\cdot\nicefrac{{\delta_{n}}}{{{\pi_{\min}}}}}}\right)+\epsilon_{n}$
$=\mathopen{}\mathclose{{\left(1-\epsilon_{n}}}\right)\cdot\mathopen{}\mathclose{{\left(2-C+\mathopen{}\mathclose{{\left(C-2}}\right)\cdot\nicefrac{{\delta_{n}}}{{{\pi_{\min}}}}}}\right)+\epsilon_{n}$
$=2-C+\mathopen{}\mathclose{{\left(C-2}}\right)\cdot\nicefrac{{\delta_{n}}}{{{\pi_{\min}}}}-\epsilon_{n}\mathopen{}\mathclose{{\left(2-C+\mathopen{}\mathclose{{\left(C-2}}\right)\cdot\nicefrac{{\delta_{n}}}{{{\pi_{\min}}}}}}\right)+\epsilon_{n}$
$=2-C+\mathopen{}\mathclose{{\left(C-2}}\right)\cdot\nicefrac{{\delta_{n}}}{{{\pi_{\min}}}}+\epsilon_{n}\mathopen{}\mathclose{{\left(C-1+\mathopen{}\mathclose{{\left(2-C}}\right)\cdot\nicefrac{{\delta_{n}}}{{{\pi_{\min}}}}}}\right)$
$=2-C+\mathopen{}\mathclose{{\left(C-2}}\right)\cdot\nicefrac{{\delta_{n}}}{{{\pi_{\min}}}}+\epsilon_{n}\,\mathopen{}\mathclose{{\left(C-1}}\right)+\epsilon_{n}\cdot\mathopen{}\mathclose{{\left(2-C}}\right)\cdot\nicefrac{{\delta_{n}}}{{{\pi_{\min}}}}$
$=2-C+\mathopen{}\mathclose{{\left(C-2}}\right)\cdot\nicefrac{{\delta_{n}}}{{{\pi_{\min}}}}+\frac{2C\delta_{n}\cdot\mathopen{}\mathclose{{\left(C-1}}\right)}{\pi_{\min}\mathopen{}\mathclose{{\left(C-1}}\right)+2C\delta_{n}}+\frac{2C\delta_{n}\cdot\mathopen{}\mathclose{{\left(2-C}}\right)}{\pi_{\min}\mathopen{}\mathclose{{\left(C-1}}\right)+2C\delta_{n}}\cdot\nicefrac{{\delta_{n}}}{{{\pi_{\min}}}}$
$=2-C+\delta_{n}\cdot\mathopen{}\mathclose{{\left(\frac{C-2}{\pi_{\min}}+\frac{2C\cdot\mathopen{}\mathclose{{\left(C-1}}\right)}{\pi_{\min}\mathopen{}\mathclose{{\left(C-1}}\right)+2C\delta_{n}}+\frac{2C\delta_{n}\cdot\pi_{\min}^{-1}\cdot\mathopen{}\mathclose{{\left(2-C}}\right)}{\pi_{\min}\mathopen{}\mathclose{{\left(C-1}}\right)+2C\delta_{n}}}}\right)$
$\geq 2-C.$
To see how we obtain the last line, note that it suffices to show the content of the parenthesis multiplied by $\delta_{n}$ is greater than zero, i.e.,
$\frac{C-2}{\pi_{\min}}+\frac{2C\cdot\mathopen{}\mathclose{{\left(C-1}}\right)}{\pi_{\min}\mathopen{}\mathclose{{\left(C-1}}\right)+2C\delta_{n}}+\frac{2C\delta_{n}\cdot\pi_{\min}^{-1}\cdot\mathopen{}\mathclose{{\left(2-C}}\right)}{\pi_{\min}\mathopen{}\mathclose{{\left(C-1}}\right)+2C\delta_{n}}$ $\geq 0$
$\iff$ $\frac{2C\cdot\mathopen{}\mathclose{{\left(C-1}}\right)+2C\delta_{n}\cdot\pi_{\min}^{-1}\cdot\mathopen{}\mathclose{{\left(2-C}}\right)}{\pi_{\min}\mathopen{}\mathclose{{\left(C-1}}\right)+2C\delta_{n}}$ $\geq\frac{2-C}{\pi_{\min}}$
$\iff$ $2C\pi_{\min}\cdot\mathopen{}\mathclose{{\left(C-1}}\right)+2C\delta_{n}\cdot\mathopen{}\mathclose{{\left(2-C}}\right)$ $\geq\mathopen{}\mathclose{{\left(2-C}}\right)\cdot\mathopen{}\mathclose{{\left(\pi_{\min}\mathopen{}\mathclose{{\left(C-1}}\right)+2C\delta_{n}}}\right)$
$\iff$ $2C\pi_{\min}\cdot\mathopen{}\mathclose{{\left(C-1}}\right)$ $\geq\mathopen{}\mathclose{{\left(2-C}}\right)\cdot\mathopen{}\mathclose{{\left(\pi_{\min}\mathopen{}\mathclose{{\left(C-1}}\right)+2C\delta_{n}-2C\delta_{n}}}\right)$
$\iff$ $2C\pi_{\min}\cdot\mathopen{}\mathclose{{\left(C-1}}\right)$ $\geq\mathopen{}\mathclose{{\left(2-C}}\right)\cdot\mathopen{}\mathclose{{\left(\pi_{\min}\mathopen{}\mathclose{{\left(C-1}}\right)}}\right)$
$\iff$ $2C$ $\geq{2-C},$
which is always satisfied because $C\geq 1$. Therefore, since this holds for any $s\in\mathcal{S}$ and both $\pi^{\prime}_{n}$ and $\pi_{n}$ have the same support, we have that $D^{\inf}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)\geq 2-C$.
Thus, we have $D^{\inf}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)\geq 2-C$ and $D^{\sup}_{\text{IR}}\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)\leq C$, $\pi^{\prime}_{n}\in\mathcal{N}^{C}\mathopen{}\mathclose{{\left(\pi_{n}}}\right)$. Therefore, $\mathcal{N}^{C}$ is lhc.
Since $\mathcal{N}^{C}$ is uhc and lhc, it is continuous. This concludes the proof of item 1.
It remains to show item 3. Let $\epsilon=\mathopen{}\mathclose{{\left(C-1}}\right)\cdot\min_{s,a}\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)$, with $a$ taken from $\text{supp}\mathopen{}\mathclose{{\left(\pi\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)}}\right)$. Assume $d\mathopen{}\mathclose{{\left(\pi,\pi^{\prime}}}\right)\leq\epsilon$ (cf.  Eq. 7). For all $s\in\mathcal{S},a\in\text{supp}\mathopen{}\mathclose{{\left(\pi\mathopen{}\mathclose{{\left({\cdotp}\mid s}}\right)}}\right)$, we have
$\pi^{\prime}\mathopen{}\mathclose{{\left(a\mid s}}\right)$
$\leq\,$ $\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)+\epsilon$
$\leq\,$ $\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)+\mathopen{}\mathclose{{\left(C-1}}\right)\cdot\min_{s,a}\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)$
$\leq\,$ $\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)+\mathopen{}\mathclose{{\left(C-1}}\right)\cdot\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)$
$=\,$ $\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)\cdot\mathopen{}\mathclose{{\left(1+C-1}}\right)$
$=\,$ $\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)\cdot C,$
or equivalently:
$\frac{\pi^{\prime}\mathopen{}\mathclose{{\left(a\mid s}}\right)}{\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)}\leq C.$
It remains to show the lower bound:
$\pi^{\prime}\mathopen{}\mathclose{{\left(a\mid s}}\right)$ $\geq\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)-\epsilon$
$=\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)-\mathopen{}\mathclose{{\left(C-1}}\right)\cdot\min_{s,a}\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)$
$\geq\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)-\mathopen{}\mathclose{{\left(C-1}}\right)\cdot\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)$
$=\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)\cdot\mathopen{}\mathclose{{\left(1-C+1}}\right)$
$=\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)\cdot C$
$\geq\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)\mathopen{}\mathclose{{\left(2-C}}\right),$
or equivalently:
$\frac{\pi^{\prime}\mathopen{}\mathclose{{\left(a\mid s}}\right)}{\pi\mathopen{}\mathclose{{\left(a\mid s}}\right)}\geq 2-C.$
This concludes the proof of item 3. ∎
Then, Theorem 1 is obtained as a corollary of Lemma 1, and the fact that the update process
$\pi_{n+1}\coloneqq\operatorname*{arg\,sup}_{\pi^{\prime}\in\mathcal{N}^{C}\mathopen{}\mathclose{{\left(\pi_{n}}}\right)}\,\mathop{\mathbb{E}}_{s\sim\xi_{\pi_{n}}}\mathop{\mathbb{E}}_{a\sim\pi^{\prime}\mathopen{}\mathclose{{\left({\cdotp}\mid{s}}}\right)}\left[A^{\pi_{n}}\mathopen{}\mathclose{{\left(s,a}}\right)\right],$
is an instance of mirror learning (Kuba et al., 2022).
Appendix C Crude Wasserstein Upper Bound
Lemma 2.
Let $s\in\mathcal{S}$ and $a\in\mathcal{A}$, the following upper bound holds:
$\mathcal{W}\big(\phi_{\sharp}P(\cdot\mid s,a),\;\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu(\cdot\mid\phi(s),a)\big)\leq\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\;\mathbb{E}_{\bar{s}^{\prime}\sim\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu(\cdot\mid\phi(s),a)}\,\,\overline{\!{d}}\big(\phi(s^{\prime}),\bar{s}^{\prime}\big).$
Proof.
$\mathcal{W}\big(\phi_{\sharp}P(\cdot\mid s,a),\;\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu(\cdot\mid\phi(s),a)\big)$
$=\sup_{\lVert f\rVert_{\text{Lip}}\leq 1}\;\Bigg[\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}f\big(\phi(s^{\prime})\big)-\mathbb{E}_{\bar{s}^{\prime}\sim\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu(\cdot\mid\phi(s),a)}f(\bar{s}^{\prime})\Bigg]$ (1)
$\leq\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\Bigg[\sup_{\lVert f\rVert_{\text{Lip}}\leq 1}\;f\big(\phi(s^{\prime})\big)-\mathbb{E}_{\bar{s}^{\prime}\sim\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu(\cdot\mid\phi(s),a)}f(\bar{s}^{\prime})\Bigg]$
$=\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\mathcal{W}\big(\delta_{\phi(s^{\prime})},\;\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu(\cdot\mid\phi(s),a)\big)$
$=\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\Bigg[\min_{\lambda\in\Lambda(\delta_{\phi(s^{\prime})},\,\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu(\cdot\mid\phi(s),a))}\mathbb{E}_{(\bar{s}_{1},\bar{s}_{2})\sim\lambda}\,d(\bar{s}_{1},\bar{s}_{2})\Bigg]$ (2)
$=\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\;\mathbb{E}_{\bar{s}^{\prime}\sim\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu(\cdot\mid\phi(s),a)}\,\,\overline{\!{d}}\big(\phi(s^{\prime}),\bar{s}^{\prime}\big).$
Here, (1) corresponds to the dual Kantorovich–Rubinstein formulation (Kantorovich and Rubinstein, 1958) where $\left\|\cdot\right\|_{\text{Lip}}$ corresponds to the Lipschitz norm, while (2) follows from the primal Monge formulation (Monge, 1781), with a trivial coupling induced by $\delta_{\phi\mathopen{}\mathclose{{\left(s^{\prime}}}\right)}$, the Dirac measure with impulse $\phi\mathopen{}\mathclose{{\left(s^{\prime}}}\right)$. ∎
Appendix D Remark on Safe Policy Improvement Methods
Standard principled safe policy improvement methods (SPI; Thomas et al., 2015; Ghavamzadeh et al., 2016a; Laroche et al., 2019; Simão et al., 2020; Castellini et al., 2023; Wienhöft et al., 2023) do not consider representation learning. Instead, SPI methods assume $\overline{\!{\mathcal{\mathcal{S}}}}\coloneqq\mathcal{S}$ and learn $\overline{\!{{R}}}$, $\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu$ by maximum likelihood estimation with respect to the experience stored in $\mathcal{B}$ collected by the behavioral $\pi_{\text{b}}$. Then, the policy improvement relies on finding the best policy in $\overline{\!{\mathcal{M}}}$ that is (probably approximately correctly) guaranteed to improves on the behavioral policy (up to an error term $\zeta>0$) against a set of all admissible MDPs, called robust MDPs (Iyengar, 2005; Nilim and Ghaoui, 2005; Wiesemann et al., 2013; Ghavamzadeh et al., 2016b; Suilen et al., 2024):
$\operatorname*{arg\,sup}_{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\in\mkern 1.5mu\overline{\mkern-1.5mu\Pi\mkern-1.5mu}\mkern 1.5mu}\rho\mathopen{}\mathclose{{\left(\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu,\,\overline{\!{\mathcal{M}}}}}\right)$ such that $\operatorname*{arg\,inf}_{\mathcal{M}^{\prime}\in\Xi\mathopen{}\mathclose{{\left(\,\overline{\!{\mathcal{M}}},e}}\right)}\rho\mathopen{}\mathclose{{\left(\pi,\mathcal{M}^{\prime}}}\right)\geq\rho\mathopen{}\mathclose{{\left(\pi_{\text{b}},\mathcal{M}^{\prime}}}\right)-\zeta,\text{ where}$
$\Xi\mathopen{}\mathclose{{\left(\,\overline{\!{\mathcal{M}}},e}}\right)\coloneqq\left\{\mathcal{M}=\langle\mathcal{S},\mathcal{A},P,{R},s_{\mathit{I}},\gamma\rangle\,\bigg|\!\begin{array}[]{rl}\left|{R}\mathopen{}\mathclose{{\left(s,a}}\right)-\,\overline{\!{{R}}}\mathopen{}\mathclose{{\left(s,a}}\right)\right|\!\!\!\!&\leq{R}_{\textsc{max}}\cdot e\mathopen{}\mathclose{{\left(s,a}}\right)\quad\quad\text{ and}\\[2.84526pt] d_{{TV}}\mathopen{}\mathclose{{\left(P\mathopen{}\mathclose{{\left({\cdotp}\mid s,a}}\right),\,\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left({\cdotp}\mid s,a}}\right)}}\right)\!\!\!\!&\leq e\mathopen{}\mathclose{{\left(s,a}}\right)\quad\quad\forall s\in\mathcal{S},a\in\mathcal{A}\end{array}\right\},$
$e\mathopen{}\mathclose{{\left(s,a}}\right)$ being an error term depending on the number of times each state $s$ and action $a$ are present in the dataset $\mathcal{B}$, and $d_{{TV}}$ being the total variation distance (Müller, 1997) which boils down to the $L_{1}$ distance when the state-action space is finite. To provide probably approximately correct (PAC) guarantees, the state-action pairs need to be visited a sufficient amount of time, depending on the size of the state-action space, to ensure $e$ is sufficiently small.
Note that the reward and total variation constraints are very related to our local losses $L_{{R}}$ and $L_{P}$: the representation corresponds here to the identity and $d_{{TV}}$ coincides with Wasserstein as the state space is discrete (Villani, 2009). The major difference here is that the bounds need to hold globally, i.e., for all state-action pairs, which make their computation typically intractable in complex settings (e.g., high-dimensional feature spaces).
We argue this objective is ill-suited to complex settings. First, classic SPI does not apply to general spaces. Second, assuming we deal with finite, high-dimensional feature spaces (e.g., visual inputs or the RAM of a video game), it is simply unlikely that $\mathcal{B}$ contains all state-action pairs. SPI with baseline bootstrapping (Laroche et al., 2019) allows bypassing this requirement by updating $\pi_{\text{b}}$ only in state-action pairs where a sufficient number of samples are present in $\mathcal{B}$. Nevertheless, this number is gigantic and is linear in the state-action space while being exponential in the size of the encoding of $\gamma$ and the desired error $\zeta$. This deems the policy update intractable. Finally, as mentioned, standard SPI does not consider representation learning. This is a further obstacle to its application in complex settings.
Appendix E Safe Policy Improvements: Proofs
Notations Henceforth, we denote by